OpenStudy (anonymous):

Which ratio is not equivalent to 45/120?

OpenStudy (anonymous):

a. 3/8 b.. 5/12 c. 9/24 d. 90/240

OpenStudy (uh60blackhawk):

will one of the easiest way to solve this problem is make every fraction into decimals

OpenStudy (uh60blackhawk):

sorry for the bad grammar it is 2 in the morning.

OpenStudy (anonymous):

ok so the decimals would be 45/120= .375 3/8= .375 9/24= .375 5/12= 0.416667 90/240=.375 thanks so its b, i appreciate u helping me!

OpenStudy (uh60blackhawk):

won't it be c because it is not .375

OpenStudy (uh60blackhawk):

oh nvm you have it in different order.

OpenStudy (anonymous):

yea i just noticed that too, thx for showing a easy strategy

OpenStudy (uh60blackhawk):

your welcome just make sure you know how to simplify fractions too. I had a teacher that won't let us use a calculator.